# STAMP/STPA informed characterization of Factors Leading to Loss of Control in AI Systems

### v1

@@ 2.1.2 Literature
Active vs passive loss of control (Bengio, 2025): In active loss of control the AI system behaves in ways that actively undermine human control. In passive loss of control people stop exercising meaningful oversight over AI systems for a variety of reasons
@@
Other literature focuses on how to analyze loss of control risk
@@ 4 Overview of Functional Control System Archetypes
Figure 4: Interacting control systems, hierarchical (vertical), and adjacent(horizontal).
@@
Figure 6: Control systems affected by control-subverting agents or actors.
@@ 4.1 Scope of Presented Research: Control System Archetype Selection
The scope of detailed characterization considered in the remainder of this paper is shown in Figure 7, where we also narrow scope by focusing on the operating life-cycle phase (not the design, development or other lifecycle phases)
Figure 7: Control system archetype considered in this paper.
@@ 5 Overview of Proposed Loss of Control Characterization Framework
For systems that include AI, in Stage 3 and 4 of the STPA process (see Figure 8) the safety practitioner makes use of our Table (see Table 1 below) in order to assist them in identifying the potential causal factors leading to loss of control resulting from the inclusion of AI in their system. Figure 9 shows the causal factor characterisation visualized in diagrammatic form. The breakdown provided in Figure 9 is derived based on consideration of Figure 8 (which itself was derived from figures G.1 and Figure 2.17 of (Leveson and Thomas, 2018)). Note that in some scenarios this diagrammatic representation of the causal pathway might just show the last part of a longer causal pathway, given that the control system is a loop.
@@
An updated method for loss scenario characterization from the developers of STPA suggest that the hazard analysis strategy is better conceptualised top-down, from hazard to causal event. One reason for undertaking the analysis in this way is to avoid the preconception that the causal chain may flow forward in time, since sometimes the causal event and hazard can occur at the same time. The loss of control characterization framework may also be provided as a table, as shown in Table LABEL:tab:cause_effect_multi. The key to the column headers is provided in appendix A.
@@ 5. Control algorithm is too slow
If an AI system is able to process information and act significantly faster than the human controller, then this speed mismatch can create an asymmetry that can erode effective control. If the required feedback is or would be received too fast for the controller to reactively respond, then the controller may need to act predictively. However, prediction by the controller of a complex AI system, with high processing speed may be poor. Where the controlled process is acting adversarially, which might potentially occur without the controller’s knowledge, concepts like relative OODA (Observe, Orient, Decide and Act loop) abilities become important in determining whether controller or controlled process prevails in determining whether the system operates within its safety constraints.
@@ 1. Process model is inadequate because objectives are outer misaligned
The controller believes that the AI has been optimized for X but the reward function is actually X’ due to challenges in accurately specifying X when training an AI. Tests do not distinguish from X or X’. This leads to the controller assuming X=X’ but that this results in failure in certain conditions, as the AI pursues goal X’ in the environment where X is the desired and expected behavior, which can lead to a hazard.
@@ 6.4 Actuators, Sensors, and Delays
  • In extreme cases implement emergency shutdown of the AI or the system down in which the AI is embedded.
@@ 7 Graduated Control System Degradations Leading to Loss of Control
A control system may cease to be adequate, even though a failure (loss of control) has not yet resulted. As Leveson reports (Leveson, 2012), frequently control systems degrade over time, for example, as organizations become complacent due to the absence of failures having occurred, or when process models in the minds of human controllers become outdated and no longer accurately reflect the a changed controlled process. At any one time, a control system might therefore contain a number of such ‘vulnerabilities’, and it is often only a matter of time before some triggering event actually causes these vulnerabilities to be manifested and for the system to fail. The loss of control event frequently includes the ‘exploitation’ of multiple such vulnerabilities. According to Leveson (Leveson, 2012), an incorrect response in the resulting accident analysis, is to focus on mitigating the particular event that caused the failure to occur. Rather the focus should be on the organizational aspects that led to the inadequate control system design, implementation or operation, or which allowed the adequacy of the control system to degrade. To mitigate such degradations in a control system, processes need to be put in place, such as performance audits that regularly check for degradations and vulnerabilities, and management of change procedures aimed at preventing degradations and vulnerabilities being introduced when system changes are implemented. These signals from audits and change management procedures (or the lack of such signals, due to the lack of audits and procedures) are the weak signals that can be detected prior to a loss of control incident actually occurring.
@@ 11 Glossary
Hazard A system state or set of conditions that, together with a particular set of worst-case environmental conditions, will lead to a loss. (Leveson and Thomas, 2018) The three basic criteria for defining system-level hazards: system states or conditions; will lead to a loss in some worst-case environment; must describe states or conditions to be prevented.
@@ Appendix B Example Application of Framework: National Intelligence Agency Use of AI Agents for Chat Monitoring
In this section we apply the framework to an example case, demonstrating its use to draw out specific loss of control scenarios that may arise. The case considers causal factors leading to loss of control in an intelligence chat monitoring system. It covers sample scenarios related to each of the control system components discussed in sections 6.2 – 6.5. While these are presented in this way, it is notable that many of the scenarios involve multiple components and interactions between them leading to the loss. (For example, sensor scenario I.a. in Table 4 involves autonomous action by the AI agent, the controller having a flawed process model, and deception via the sensor.)

### v2

@@ 2.1.2 Literature
Active vs passive loss of control (Bengio, 2025): In active loss of control the AI system behaves in ways that actively undermine human control. In passive loss of control people stop exercising meaningful oversight over AI systems for a variety of reasons.
@@
Other literature focuses on how to analyze loss of control risk.
@@ 4 Overview of Functional Control System Archetypes
Figure 4: Interacting control systems, hierarchical (vertical), and adjacent(horizontal)
@@
Figure 6: Control systems affected by control-subverting agents or actors
@@ 4.1 Scope of Presented Research: Control System Archetype Selection
The scope of detailed characterization considered in the remainder of this paper is shown in Figure 7, where we also narrow scope by focusing on the operating life-cycle phase (not the design, development or other lifecycle phases).
Figure 7: Control system archetype considered in this paper
@@ 5 Overview of Proposed Loss of Control Characterization Framework
For systems that include AI, in Stage 3 and 4 of the STPA process (see Figure 8) the safety practitioner makes use of our Table (see Table 1 below) in order to assist them in identifying the potential causal factors leading to loss of control resulting from the inclusion of AI in their system. Figure 9 shows the causal factor characterisation visualized in diagrammatic form. The breakdown provided in Figure 9 is derived based on consideration of Figure 8 (which itself was derived from Figure G.1 and Figure 2.17 of (Leveson and Thomas, 2018)). Note that in some scenarios this diagrammatic representation of the causal pathway might just show the last part of a longer causal pathway, given that the control system is a loop.
@@
An updated method for loss scenario characterization from the developers of STPA suggest that the hazard analysis strategy is better conceptualised top-down, from hazard to causal event. One reason for undertaking the analysis in this way is to avoid the preconception that the causal chain may flow forward in time, since sometimes the causal event and hazard can occur at the same time. The full loss of control characterization table is provided in Table LABEL:tab:cause_effect_multi, Appendix A. The key to the column headers is provided in Appendix A.
@@ 5. Control algorithm is too slow
If an AI system is able to process information and act significantly faster than the human controller, then this speed mismatch can create an asymmetry that can erode effective control (Aguirre, 2025). If the required feedback is or would be received too fast for the controller to reactively respond, then the controller may need to act predictively. However, prediction by the controller of a complex AI system, with high processing speed may be poor. Where the controlled process is acting adversarially, which might potentially occur without the controller’s knowledge, concepts like relative OODA (Observe, Orient, Decide and Act loop) abilities become important in determining whether controller or controlled process prevails in determining whether the system operates within its safety constraints.
@@ 1. Process model is inadequate because objectives are outer misaligned
The controller believes that the AI has been optimized for X but the reward function is actually X’ due to challenges in accurately specifying X when training an AI. Tests do not distinguish from X or X’. This leads to the controller assuming X=X’ but this results in failure in certain conditions, as the AI pursues goal X’ in the environment where X is the desired and expected behavior, which can lead to a hazard.
@@ 6.4 Actuators, Sensors, and Delays
  • In extreme cases implement emergency shutdown of the AI or the system in which the AI is embedded.
@@ 7 Graduated Control System Degradations Leading to Loss of Control
A control system may cease to be adequate, even though a failure (loss of control) has not yet resulted. As Leveson reports (Leveson, 2012), frequently control systems degrade over time, for example, as organizations become complacent due to the absence of failures having occurred, or when process models in the minds of human controllers become outdated and no longer accurately reflect a changed controlled process. At any one time, a control system might therefore contain a number of such ‘vulnerabilities’, and it is often only a matter of time before some triggering event actually causes these vulnerabilities to be manifested and for the system to fail. The loss of control event frequently includes the ‘exploitation’ of multiple such vulnerabilities. According to Leveson (Leveson, 2012), an incorrect response in the resulting accident analysis, is to focus on mitigating the particular event that caused the failure to occur. Rather the focus should be on the organizational aspects that led to the inadequate control system design, implementation or operation, or which allowed the adequacy of the control system to degrade. To mitigate such degradations in a control system, processes need to be put in place, such as performance audits that regularly check for degradations and vulnerabilities, and management of change procedures aimed at preventing degradations and vulnerabilities being introduced when system changes are implemented. These signals from audits and change management procedures (or the lack of such signals, due to the lack of audits and procedures) are the weak signals that can be detected prior to a loss of control incident actually occurring.
@@ 11 Glossary
Hazard A system state or set of conditions that, together with a particular set of worst-case environmental conditions, will lead to a loss (Leveson and Thomas, 2018). The three basic criteria for defining system-level hazards: system states or conditions; will lead to a loss in some worst-case environment; must describe states or conditions to be prevented.
@@ Appendix B Example Application of Framework: National Intelligence Agency Use of AI Agents for Chat Monitoring
In this section we apply the framework to an example case, demonstrating its use to draw out specific loss of control scenarios that may arise. The case considers causal factors leading to loss of control in an intelligence chat monitoring system. It covers sample scenarios related to each of the control system components discussed in sections 6.1 – 6.4. While these are presented in this way, it is notable that many of the scenarios involve multiple components and interactions between them leading to the loss. (For example, the sensor scenario in Table 4 involves autonomous action by the AI agent, the controller having a flawed process model, and deception via the sensor.)